OpenStudy (anonymous):

F=(x) 4x^3 - 12x^2 -16x WHat are the zeros of the function? What are their multiplicties?.

OpenStudy (anonymous):

The answer is the numbers -1, 4, and 0 are zeros of the multiplicity.

OpenStudy (anonymous):

Hoow?

OpenStudy (anonymous):

\[ 4x^3 - 12x^2 -16x\] \[=4x(x^2-3x-4)\] is a start

OpenStudy (anonymous):

then factor as \[4x(x-4)(x+1)\]

OpenStudy (anonymous):

Oh factoring lmao =.=

OpenStudy (anonymous):

sorry for the bother

OpenStudy (anonymous):

\(4x=0\iff x=0. x-4=0\iff x=4,x+1=0\iff x=-1\)